Jeff takes 5 hr longer to build a fence than it takes Bill. When they work together, it takes them 6 hours? How long would it take Bill to do the job alone?
Answer by scott8148(6628)

let x="Bill's time", so x+5="Jeff's time"

if it takes h hours to complete a task; then in one hour, 1/h is the fraction of the task completed
___ all of the fractions sum to 1, the whole task

(6 / x) + (6 / (x+5)) = 1

multiplying by x(x+5) ___ 6x + 30 + 6x = x^2 + 5x ___ 0 = x^2 - 7x - 30

factoring ___ 0 = (x - 10)(x + 3)

so, x = 10 (Bill)